OpenStudy (anonymous):

simplify the expression (3x+5)(x-4)/(3x-5)(4-x)

OpenStudy (anonymous):

do you know how to FOIL?

OpenStudy (anonymous):

no

OpenStudy (anonymous):



OpenStudy (anonymous):

the arrows are what you multiply and 1-4 is the order most people use

hero (hero):

\[\frac{(3x+5)(x-4)}{(3x-5)(4-x)}\] \[-\frac{(3x+5)(x-4)}{(3x-5)(x-4)}\] \[-\frac{(3x+5)}{(3x-5)}\]

OpenStudy (anonymous):

ok